OpenStudy (trwatkins1):

–12 – 4.8 ∙ (–2 + –8) A. –168 B. 36 C. 72 D. 168

OpenStudy (trwatkins1):

that was quick

OpenStudy (youngstudier):

XD

OpenStudy (youngstudier):

What do u think the answer is?

OpenStudy (trwatkins1):

u readt

OpenStudy (trwatkins1):

*ready

OpenStudy (youngstudier):

?

OpenStudy (youngstudier):

Yep, whatcha think the answer is?

OpenStudy (trwatkins1):

36?

OpenStudy (trwatkins1):

thats what i thnk the answer is

OpenStudy (anonymous):

Break the problem down, start with this: -12-4.8=x then -2+(-8)=y then x+y=Answer

OpenStudy (youngstudier):

I believe B is correct, my thing was glitching again.

OpenStudy (anonymous):

mine too

OpenStudy (anonymous):

I would say it's A...